OpenStudy (anonymous):

I have a homework problem that I could use some help with if anyone knows the answer. "A ball is thrown upward with a speed of 22m/s. What is the ball's height after 2s"

OpenStudy (lalaly):

S = VoT + (1/2)gT^2 where S = ball's attained height Vo = 22 m/sec. T = 2 sec. g= 9.8 m/sec^2 Substituting S = 22(2) + (1/2)(-9.8)(2^2)

OpenStudy (lalaly):

the negative sign (-9.8 m/sec^2). This simply implies that the ball is being slowed down by gravity as it is going up.

OpenStudy (anonymous):

Cool. Thanks a lot!